There are 3 mixtures when mixed in the ratio of 3:4:5 price was 240.When mixed with 6:4:9 price was 340.what's the price when mixed in 5:6:8 ration price is_________

Grade:11

1 Answers

Aman Bansal
13 years ago

 dear doppalapudi,

the mixtures are x, y and z
3x + 4y + 5z = 240
6x + 4y + 9z = 340
eliminate y
3x + 4z = 100
3x = 100 - 4z
3x = 4*25 - 4z
assuming x, y and z are integers, x must be divisible by 4
let x = 4A where A is an integer
3*4A = 4*25 - 4z
3A = 25 - z
25 - z must be divisible by 3, possible values for z are
1, 4, 7, 10, 13, 16, 19, 22

from
6x + 4y + 9z = 340
3x + 4y + 5z = 240
multiply by 2
6x + 8y + 10z = 480
eliminate x
4y + z = 140
z = 140 - 4y
z = 4*35 - 4y
z must be divisible by 4, thus
z = 4 or 16
possible solutions for (x,y,z) are
(28,34,4) and (12,31,16)

with a ratio of 5:6:8,
5*28 + 6*34 + 8*4 = 376
5*12 + 6*31 + 8*16 = 374

We are all IITians and here to help you in your IIT JEE preparation.

Now you can win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar : Click here to download the toolbar.. 

AMAN BANSAL